OpenStudy (anonymous):

I am needing help if someone can explain how to find the solution set for (x^2-6x)^2-35(x^2-6x)-200=0

OpenStudy (anonymous):

solve \[z^2-35z-200=0\] the replace z by \[x^2-6x\] and solve those

OpenStudy (anonymous):

solutions to first one is ok because it factors as \[(z-40)(z+5)=0\] \[z=40,z=-5\] then solve \[x^2-6x=40\] and \[x^2-6x=5\] to finish up

OpenStudy (anonymous):

So as I what was mentioned above I get x=10, -4, 5, 1????